OpenStudy (anonymous):

What are the solutions of the equation? c^2-10c=0 A. 1, -√10 B. 0, –10 C. 0, √10 D. 0, 10 A, B, C, or D? Please explain :)

jimthompson5910 (jim_thompson5910):

c^2-10c=0 c(c-10) = 0 See where to go next?

OpenStudy (anonymous):

No, I do not.

jimthompson5910 (jim_thompson5910):

c(c-10) = 0 breaks up to c = 0 or c-10 = 0

jimthompson5910 (jim_thompson5910):

then what

OpenStudy (anonymous):

I don't know. I'm not good at algebra.

jimthompson5910 (jim_thompson5910):

solve c-10 = 0 for c

OpenStudy (anonymous):

I don't know how.

jimthompson5910 (jim_thompson5910):

something - 10 = 0

jimthompson5910 (jim_thompson5910):

what number goes in the place of "something"

OpenStudy (anonymous):

Important step: If \[ a\cdot b=0 \]Then at least one of \(a, b\) must be zero.

OpenStudy (anonymous):

10 - 10 = 0 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so c-10 = 0 means that c = 10

jimthompson5910 (jim_thompson5910):

so c = 0 or c = 10

OpenStudy (anonymous):

So the answer is D?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Ah, okay thank you :)